# Uncovering thermodynamic origin of counterflow and coflow instabilities in miscible binary superfluids

### v1

@@ Abstract
In this paper, we explore instabilities in binary superfluids with a nonvanishing relative superflow, particularly focusing on counterflow and coflow instabilities. We extend recent results on the thermodynamic origin of finite superflow instabilities in single-component superfluids to binary systems and derive a criterion for the onset of instability through a hydrodynamic analysis. To verify this result, we utilize both the Gross-Pitaevskii equation (GPE) for weakly interacting Bose-Einstein condensates (BEC) and a holographic binary superfluid model, which naturally incorporates strong coupling, finite temperature, and dissipation. We find that the counterflow and coflow instabilities in binary superfluids are all essentially thermodynamic. Except the one due to order competing via global thermodynamic instability, the others are caused by an eigenvalue of the free energy Hessian diverging and changing sign. We also observe that the critical velocities of these instabilities follow a general scaling law related to the interaction strength between superfluid components. The nonlinear stages of the instabilities are also studied by full time evolution, where vortex dynamics is found to play a significant role, resulting in the reduction of superfluid velocity back to a stable phase.
I Introduction
@@
In both classical fluid systems and quantum superfluids with nonzero velocity, instabilities, such as the Kelvin-Helmholtz instability Helmholtz (1868); Thomson (1871); Takeuchi et al. (2010a); Volovik (2002); Blaauwgeers et al. (2002); Finne et al. (2006); Eltsov et al. (2019), and the two-stream instability Abad et al. (2015); Law et al. (2001); Ishino et al. (2011); Takeuchi et al. (2010b); Schmitt (2014); Delehaye et al. (2015); Andersson et al. (2004); Haber et al. (2016), appear very generally. Such instabilities typically develop into turbulence in nonlinear stage. Understanding the origin and the final fate of these instabilities is crucial to understanding turbulence. Binary superfluids are the most suitable platform to study these instabilities since there is no friction between the two components and one can study equilibrium states with nonzero relative velocity. Such an instability in miscible binary superfluids is often called counterflow instability. The counterflow instability of weakly interacting Bose-Einstein condensates (BECs) has been extensively studied Abad et al. (2015); Law et al. (2001); Ishino et al. (2011); Takeuchi et al. (2010b). However, these studies mostly rely on Gross-Pitaevskii equation (GPE), which is only valid in weak coupling limit, zero temperature, and does not include dissipation.
On the other hand, Gauge-Gravity duality, also known as holography or AdS/CFT (Anti-de Sitter/Conformal Field Theory), has been widely used to study strongly-correlated condensed matter systems without quasiparticles (see Hartnoll et al. ; Zaanen et al. (2015) for comprehensive reviews). This method connects a pure classical gravity theory to a strong coupling quantum field theory without dynamical gravity on the boundary of the spacetime. Finite temperature, finite density states are modeled by a charged black hole in the bulk. Therefore, in order to study instabilities in binary superfluids, in this work we go beyond the regime of validity of GPE by employing a holographic binary superfluid model, which is intrinsically in the strong coupling limit, and naturally incorporates finite temperature and dissipation. In previous works by some of us An et al. (2024a, b), we explored dynamical interface instability of strongly interacting binary superfluids. This kind of instability occurs at the interface of immiscible binary superfluids with nonzero superfluid velocity, corresponding to Kelvin-Helmholtz instability. In this work, we focus on instabilities in miscible binary superfluids.
More recently, one of us argued that the dynamical instability under linear perturbations of interacting systems in the hydrodynamic regime generally follows from local thermodynamic instability together with positively of entropy production Goutéraux et al. (2023); Areán et al. (2024); Goutéraux and Mefford (2024). In particular, the Landau instability in single component superfluids was shown to be essentially thermodynamic. In this work, we extend this result to binary superfluids. By resorting to hydrodynamics of homogeneous binary superfluids, we determine the criterion for thermodynamic instability. To verify this criterion, we analyse both the weakly interacting GPE, and the strongly interacting holographic model. We find that instabilities in both models are essentially thermodynamic.
For GPE, the critical velocity for counterflow instability has already been known for some time Abad et al. (2015); Law et al. (2001); Ishino et al. (2011); Takeuchi et al. (2010b). In this work we show that the critical instability is correctly predcited by the criterion for thermodynamic instability. For the holographic model, by analysing the spectrum of quasinormal modes (QNMs) on top of stationary states in this model, we identify various instabilities and their critical velocities. Beside the usual counterflow instability, we also find a coflow instability in the holographic model, when the two components flow with the same velocity in the same direction. These instabilities are also identified to be thermodynamic by the same criterion. Interestingly, beyond the criterion from hydrodynamic analysis, there is a global thermodynamic instability due to order competing.
@@
Further, both in GPE and in the holographic model, we find that the critical velocities of these instabilities present a general scaling law with respect to interaction strength between the two components.
@@
This paper is organized as follows. In section II, we present the theory of homogeneous binary superfluids and establish the instability criterion, see equation (17) below. In section III, we study the counterflow instability with GPE and check its onset is predicted by (17). In section IV, we turn to holographic binary superfluids, study their linear instabilities and match to the criterion (17), and determine their nonlinear time evolution. We discuss the universal scaling law of the critical velocity in section V and finish with some conclusions VI. Finally, more details and discussions appear in the appendices.
II Hydrodynamics of homogeneous binary superfluids
In this section, we extend the formulation in Goutéraux et al. (2023); Areán et al. (2024); Goutéraux and Mefford (2024) to binary superfluids for which there can be in general two conserved $U(1)$ charges. Then the conservation equations read  [^1] While we find it convenient to write the conservation equations and constitutive relations for the spatial fluxes non-covariantly, for the purposes of this work we always have in mind systems ultimately invariant under Galilean or Lorentz boost, which set $g^{i}=j^{i}_{1}+j^{i}_{2}$ and $g^{i}=j_{\epsilon}^{i}$, respectively.
@@ II Hydrodynamics of homogeneous binary superfluids
In this paper we work in the grand-canonical ensemble where the thermodynamic variables are the temperature $T$, the chemical potential $\mu$ and the norm of the superfluid velocities $v_{I}^{s}=|\bm{v^{s}_{I}}|$. The first law of thermodynamics reads
$d\epsilon=$ $Tds+\mu^{I}dn_{I}+\bm{v^{n}}d\bm{g}$ (3)
$+\bm{h^{I}}d\bm{v^{s}_{I}}+Nd(\bm{v^{s}_{1}}\cdot\bm{v^{s}_{2}})\,,$
where $s$ is the entropy density, $\bm{v^{n}}$ is the normal fluid velocity, and $\bm{h_{I}}=n^{s}_{I}(\bm{v^{s}_{I}}-\bm{v^{n}})$ with $n^{s}_{I}$ being the superfluid charge density ($I$ is not summed here). In contrast to the single component case Goutéraux et al. (2023), there is an additional cross term which accounts for the energy change due to relative motion between the two components of superfluids. [^2] In principle in absence of boost invariance there is a relative coefficient between $\bm{v^{s}_{I}}$ and $\bm{v^{n}}$ in $\bm{h_{I}}$ Armas and Have (2024). As it plays no role in our analysis, we have set it to unity. By introducing $\bm{\bar{h}_{I}}=\bm{h_{I}}+N\sigma^{IJ}\bm{v^{s}_{J}}$ with $\sigma^{IJ}=\left(\begin{smallmatrix}0&1\\ 1&0\end{smallmatrix}\right)$, this cross term can be absorbed and one has
@@
We now consider dissipative corrections to the idea order constitutive relations, for which we shall use tilde to label dissipative corrections. Then from positivity of entropy production
$T\partial_{t}s+T\partial_{i}(sv^{ni}+\tilde{j}^{i}_{s}/T)\equiv\Delta\geq 0\,,$ (5)
we can obtain the constitutive relations for the currents in equilibrium state compatible with parity and time-reversal invariance:
@@
where $p=-\epsilon+sT+n_{I}\mu_{I}+v^{ni}g_{i}$, and tilded quantities are first-order derivative corrections. To first order, the entropy production is given by
$\Delta=-\tilde{j}_{s}^{i}\partial_{i}T/T-\tilde{j}_{I}^{i}\partial_{i}\mu^{I}-\tilde{\tau}^{ji}\partial_{i}v^{n}_{j}-\tilde{\mu^{I}}\partial_{i}\bar{h}_{I}^{j}\,,$ (7)
where $\tilde{\mu^{I}}$ are derivative corrections in the Josephson equations.
Next, we turn to the dynamical instability from hydrodynamics. One can linearize the equations of motion (EoMs) around equilibrium state by perturbing the thermodynamic quantities $O_{A}=O_{A0}+e^{-i\omega t+i\bm{k}\cdot\bm{x}}\delta O_{A}$ and their sources $s_{A}=s_{A0}+e^{-i\omega t+i\bm{k}\cdot\bm{x}}\delta s_{A}$, where $O_{A}=(s,\bm{g},n_{I},\bm{v^{s}_{I}})$ and $s_{A}=(T,\bm{v^{n}},\mu_{I},\bm{\bar{h}_{I}})$. They are related by static susceptibility matrix $\chi_{AB}=\delta O_{A}/\delta s_{B}=\delta^{2}F/\delta s_{A}\delta s_{B}$, where $F=-T\mathrm{ln}Z$ is the thermal free energy. The linearized EoMs in Fourier space would be

### v2

# Counterflow and coflow instabilities in miscible binary superfluids

@@ Abstract
We explore instabilities in binary superfluids with a nonvanishing relative superflow, particularly focusing on counterflow and coflow instabilities. We extend recent results on the thermodynamic origin of finite superflow instabilities in single-component superfluids to binary systems and derive a criterion for the onset of instability through a hydrodynamic analysis, which applies to interacting many-body systems at finite temperature. We find that the onset of these instabilities is signaled by the determinant of the Hessian of the thermal free energy diverging and changing sign. We verify this hydrodynamic prediction in a holographic binary superfluid modeled with gauge/gravity duality, which naturally incorporates strong coupling, finite temperature, and dissipation. We also compare to results obtained using the Gross-Pitaevskii equation for weakly interacting Bose-Einstein condensates and find that the same criterion continues to apply at zero temperature, where it reduces to evaluating derivatives of the supercurrents with respect to the superfluid velocities. We observe that the critical velocities of these instabilities follow a general scaling law related to the interaction strength between superfluid components. Finally, the nonlinear stages of the instabilities are studied by full time evolution using gauge/gravity duality, where vortex annihilation leads to a decrease of superfluid velocity back to a value where the binary superfluid phase is stable.
I Introduction and summary of results
@@
In both classical fluid systems and quantum superfluids with nonzero velocity, instabilities, such as the Kelvin-Helmholtz instability Helmholtz (1868); Thomson (1871); Takeuchi et al. (2010a); Volovik (2002); Blaauwgeers et al. (2002); Finne et al. (2006); Eltsov et al. (2019), and the two-stream instability Abad et al. (2015); Law et al. (2001); Ishino et al. (2011); Takeuchi et al. (2010b); Schmitt (2014); Delehaye et al. (2015); Andersson et al. (2004); Haber et al. (2016), appear very generally. Such instabilities typically develop nonlinearly into turbulence. Understanding the origin and final fate of these instabilities is crucial to understanding turbulence. Binary superfluids are an appealing platform to study these instabilities since there is no friction between the two components and one can study equilibrium states with nonzero relative velocity. This is called a counterflow instability. The counterflow instability of weakly interacting Bose-Einstein condensates (BECs) has been extensively studied theoretically, e.g. Khalatnikov (1973); Nepomnyashchii (1974, 1976); Yukalov (1980); Abad et al. (2015); Law et al. (2001); Ishino et al. (2011); Takeuchi et al. (2010b), and experimentally, e.g. Hamner et al. (2011); Beattie et al. (2013a); Kim et al. (2017), see e.g. Pethick and Smith (2008); Pitaevskii and Stringari (2003) for textbooks on this topic. In these studies, the Gross-Pitaevskii equation (GPE) plays a prominent role, and describes the dynamics of a gas of weakly-interacting bosons at zero temperature.
In this work, we study the coflow and couterflow instabilities of dissipative binary superfluids at finite temperature employing hydrodynamics and gauge/gravity duality. These approaches probe a regime different from that of the GPE. Namely, they are both applicable in the presence of strong interactions, and so do not assume the existence of underlying long-lived quasiparticles. They also allow to model thermal dissipative effects. We also comment on how our results relate to the analyses using the GPE.
I.1 Non-quasiparticle approaches to superfluids
Hydrodynamics Landau and Lifshitz (1987); Chaikin and Lubensky (1995); Kovtun (2012); Forster (2018), which we review in greater detail in Appendix A, is a celebrated and appealing framework for capturing the late-time collective dynamics of interacting systems at finite temperature, such as electron flow in Graphene Lucas and Fong (2018); Fritz and Scaffidi (2024) or the Quark-Gluon plasma formed in heavy ion collisions Luzum and Romatschke (2008). The late-time and finite temperature assumptions are crucial as they allow us to postulate the existence of a local thermodynamic equilibrium state, which then relaxes to global equilibrium. Since microscopic details of the system enter only through the thermodynamic equation of state and the so-called transport coefficients characterizing out-of-equilibrium dynamics, hydrodynamics allows one to derive universal results applying to widely different microscopic systems, which is the philosophy we adopt in this work. Specifically, the results we derive using hydrodynamics will be seen to correctly reproduce the onset of instabilities using either the GPE or gauge-gravity duality.
Gauge-gravity duality Ammon and Erdmenger (2015); Zaanen et al. (2015); Hartnoll et al. is a microscopic framework which maps the dynamics of a certain class of strongly-coupled systems (specifically, strongly-coupled non-Abelian gauge theories with an infinite-dimensional gauge group) to gravity theories defined on spacetimes with a negative global Ricci curvature, called anti-de Sitter spacetimes. See Appendix B for more details. Originally, this duality was discovered Maldacena (1998); Witten (1998); Gubser et al. (1998) in the context of String Theory, a theory of quantum gravity. Charged black holes in the gravity theory map to thermal, finite density states in the dual field theory. The area of the event horizon is dual to the entropy of the state, and its temperature is given by the Hawking temperature of the black hole. The late-time dynamics of these black holes is found to be governed by hydrodynamics, as expected on general grounds for interacting systems (whether strongly or weakly coupled). Gauge-gravity duality then offers a solvable framework to investigate and confirm hydrodynamic predictions in a class of models which do not rely on the assumption of weakly-coupled quasiparticles, and is complementary to such approaches. While the gravitational model we use in this work could in principle be connected to a precise microscopic theory by embedding it into String Theory, it suffices for our purposes that it matches hydrodynamic predictions in the infrared, irrespective of specific microscopic UV completions. In fact, this serves to emphasize the main point of this work, that the onset of the instabilities we study may be understood without referring to a specific microscopic model.
Noteworthy results of the dialogue between gauge-gravity duality and hydrodynamics are the prediction of a lower bound on the shear viscosity to entropy density ratio of strongly-coupled quantum fluids, Kovtun et al. (2005), verified by the Quark-Gluon-Plasma and trapped cold atoms gases Schäfer and Teaney (2009); and on the diffusion of quantum fluids Hartnoll (2015), such as those found in cold atoms systems Brown et al. (2019) or in the strange metal phase of high $T_{c}$ cuprate superconductors Zhang et al. (2017).
Both approaches, hydrodynamics and gauge-gravity duality, can be extended to superfluids. Superfluid hydrodynamics was devised by Landau in the context of Helium 4 Landau (1941), followed by many subsequent developments (see Schmitt (2015) for a review of relativistic superfluid hydrodynamics, Chaikin and Lubensky (1995) for Galilean superfluids). In the context of cold atoms, zero-temperature superfluids are described by ideal superfluid hydrodynamics, where temperature effects, and thus dissipation, are neglected Pitaevskii and Stringari (2016); Sonin (2016). This hydrodynamic theory should be viewed as a low-energy effective theory, and can be derived more microscopically starting from the GPE. The underlying reason why this can be done is the emergence of two global symmetries in the infrared, a global U(1) symmetry corresponding to charge conservation, and the conservation of the winding of the phase of the superfluid order parameter, which in the absence of vortices gives rise to conserved supercurrents. These global symmetries constrain the effective theory in a way that reduces it to ideal, zero-temperature superfluid hydrodynamics, even though the notion of local thermodynamic equilibrium which underlies hydrodynamics as an effective theory is missing.
In gauge-gravity duality, a superfluid is modeled by constructing a black hole endowed with scalar ‘hair’, meaning it has a ‘cloud’ of a charged, complex scalar condensate hovering outside its horizon Gubser (2008); Hartnoll et al. (2008a, b). This is dual to the condensate (order parameter) in the dual field theory. Correspondingly, the late-time dynamics of these hairy black holes maps to superfluid hydrodynamics Amado et al. (2009); Herzog and Yarom (2009); Sonner and Withers (2010); Herzog et al. (2011); Bhattacharya et al. (2011); Goutéraux and Mefford (2020a, b); Arean et al. (2021); Goutéraux et al. (2023); Areán et al. (2024). In previous works by some of us An et al. (2024a, b), the dynamical Kelvin-Helmholtz instability at the interface of strongly interacting immiscible binary superfluids was also explored using gauge-gravity duality.
I.2 Thermodynamic origin of the Landau instability of single-component superfluids
In single-component superfluids, a well-known instability occurs for states with sufficiently large superflow. An elegant intuitive explanation was given by Landau, making use of the quasi-particle description of weakly-coupled Galilean superfluids such as Helium 4. The superfluid phonons are bosonic quasiparticles and their quasiparticle energy may become negative at large enough superflow, using a Galilean transformation to boost back to the superfluid rest frame Landau et al. (1980); Schmitt (2015). In Goutéraux et al. (2023); Areán et al. (2024), it was shown that the instability can equally well be characterized by the divergence and change of sign of one of the eigenvalues of the Hessian of the free energy density $f$, which occurs when
$\left.\frac{\partial(n_{s}v_{s})}{\partial v_{s}}\right|_{T,\mu}=0\,,\quad n_{s}v_{s}=\left.\frac{\partial f}{\partial v_{s}}\right|_{T,\mu}\,.$ (1)
Here, $(n_{s},v_{s})$ are the superfluid density and the norm of the superfluid velocity, respectively, which give the superfluid contribution to the charge current. This gives a concrete connection between the microscopic description of the instability (quasiparticles are unstable) and a macroscopic one: the system becomes locally thermodynamically unstable, with a precise identification of the sector where the instability occurs. Beside Helium 4, this criterion also matches the critical supercurrent condition for BCS superconductors Bardeen (1962)
This instability is also dynamical, in the sense that time-dependent plane-wave perturbations $\sim e^{-i\omega t+ikx}$ of the state grow exponentially with time, $\textrm{Im}\,\omega>0$. This is manifested by one of the collective hydrodynamic excitations moving into the upper complex frequency plane when (1) is crossed.
This macroscopic argument can easily be extended to other systems than weakly-coupled, Galilean superfluids, and indeed it correctly predicts the onset of the instability in strongly-coupled superfluids with a gravitational dual Goutéraux et al. (2023); Areán et al. (2024) as well as many other systems Goutéraux and Mefford (2024). In Goutéraux and Mefford (2024), a general recipe was given to construct hydrodynamic theories with stable linear dynamical evolution provided that the static susceptibility matrix is positive-definite and the local second law of thermodynamics holds.
One of the main results of the present work is to extend the instability criterion (1) to binary superfluids at finite temperature and check that it correctly predicts the onset of coflow and counterflow instabilities in a gauge-gravity duality model of binary superfluids.
I.3 Summary of results
The main aim of this work is to show that dynamical instabilities which occur in binary superfluids can be predicted by computing the appropriate static susceptibility. The precise criterion is given in (22) below. We now summarize our main results.
We first write the hydrodynamics of dissipative binary superfluids, based on the standard hydrodynamic theory of superfluids Chaikin and Lubensky (1995); Pitaevskii and Stringari (2016); Sonin (2016) and going beyond previous work on binary superfluids Khalatnikov (1957, 1973); Nepomnyashchii (1976); Andreev and Bashkin (1975) by including dissipative terms to first order in derivatives of the hydrodynamic fields and by relaxing the requirement of boost invariance.
Doing so allows us to demonstrate that the onset of these instabilities is predicted by one of the eigenvalues of the Hessian of the free energy diverging and changing sign provided that the local second law of thermodynamics holds, as for single-component superfluids. Conversely, this establishes that the hydrodynamic theory of binary superfluids is stable under linear perturbations provided the local thermodynamic equilibrium is stable and the divergence of the local entropy current is positive.
The exact form of criterion for instabilities is given in (22) below. In the limit where we freeze the velocity of the normal component of the superfluid as well as temperature fluctuations, this is controlled by the dependence of the superfluid densities on the superfluid velocities.
We carefully verify that this is precisely reproduced in our holographic model of binary superfluids, by analysing the spectrum of quasi-normal modes (QNMs) on top of stationary black hole states. In brief, this means solving the Maxwell’s equations coupled to a complex scalar on the black hole background spacetime. Indeed, these QNMs directly correspond to the poles of superfluid hydrodynamics in the complex frequency plane through the holographic dictionary. In the unstable region predicted by (22), we both find that the Hessian is no longer positive-definite and that the spectrum includes a collective mode with $\textrm{Im}\,\omega>0$.
The dynamics of a weakly-interacting Bose gas described by the GPE can be mapped to perfect superfluid hydrodynamics, setting the temperature to zero and neglecting the role of dissipation Pitaevskii and Stringari (2016); Sonin (2016). Since temperature and normal velocity effects are unimportant to leading order for capturing the coflow and counterflow instabilities hydrodynamics, we expect our criterion to hold as well, when modeling the dynamics using the GPE. For the GPE, the critical velocity for counterflow instability has been computed from the microscopic Bogoliubov-de Gennes theory in previous literature Abad et al. (2015); Law et al. (2001); Ishino et al. (2011); Takeuchi et al. (2010b) and we show that this expression exactly matches the one derived from perfect binary superfluid hydrodynamics when applying our criterion (22). The reason behind this is clear: since in the perfect superfluid limit, the second law is automatically satisfied, it suffices to examine the appropriate ‘thermodynamic’ derivatives. Even at strictly zero temperature, these derivatives are well defined (we can define the charge and superfluid densities and the superfluid velocity from the microscopic quantum operators), and so our criterion continues to hold.
Thus the instability criterion we establish using the collective hydrodynamic description is found to work in widely different kinds of binary superfluids: an ideal, zero temperature superfluid with weakly-coupled Bogoliubov-de Gennes quasiparticles on one hand, and strongly-coupled binary superfluids without quasiparticles modeled using gauge-gravity duality.
@@ I.3 Summary of results
Further, both in GPE and in the holographic model, we find that the critical superfluid velocities where these instabilities occur present a general scaling law with respect to the interaction strength between the two components.
@@
This paper is organized as follows. In section II.1, we present the theory of homogeneous binary superfluids and establish the instability criterion, see equation (22) below. In section III, we study the counterflow instability with GPE and check its onset is predicted by (22). In section IV, we turn to holographic binary superfluids, study their linear instabilities and check that their onset matches the criterion (22). Finally, we determine their nonlinear time evolution. We discuss the universal scaling law of the critical velocity in section V and end with some conclusions VI. Finally, more details and discussion are given in a set of appendices , such as reviews of hydrodynamics (Appendix A) and gauge/gravity duality (Appendix B), as well as details of our holographic setup (Appendix C).
II Counterflow and coflow instabilities from binary superfluid hydrodynamics
In this section, we first review how to set up the hydrodynamics of binary superfluids in subsection II.1. While we start our treatment very generally, in section II.2 we freeze the fluctuations of the normal fluid and of the temperature in order to facilitate the analysis. Indeed, our purpose is not to formulate the hydrodynamic theory in full generality, but rather to analyze how the counterflow and coflow instabilities arise within hydrodynamics, and demonstrate that they can be predicted by a simple condition on static susceptibilities, (22).
II.1 Hydrodynamics of homogeneous binary superfluids
In this subsection, we review the formulation of binary superfluid hydrodynamics. This differs from ordinary superfluid hydrodynamics (see eg Chaikin and Lubensky (1995) for a standard treatment), as there are generally two conserved $U(1)$ charges and two independent Goldstone degrees of freedom. At ideal order and with Galilean symmetry, this is well known Andreev and Bashkin (1975). Our goal is to formulate the theory for non-Galilean invariant states (in particular so we can connect to the holographic treatment, which is Lorentz-invariant), and to account for thermal dissipation, along the lines of non-ideal superfluid hydrodynamics Chaikin and Lubensky (1995).
The conservation equations take the form of generalized continuity equations  [^1] While we find it convenient to write the conservation equations and constitutive relations for the spatial fluxes non-covariantly, for the purposes of this work we always have in mind systems ultimately invariant under Galilean or Lorentz boost, which set $g^{i}=j^{i}_{1}+j^{i}_{2}$ and $g^{i}=j_{\epsilon}^{i}$, respectively. At ideal order, boost invariance does not lead to major differences, but the number of derivative corrections proliferates rapidly without boost invariance Novak et al. (2020); de Boer et al. (2020); Armas and Jain (2021).
@@ II.1 Hydrodynamics of homogeneous binary superfluids
In this paper we work in the grand-canonical ensemble where the thermodynamic variables are the temperature $T$, the chemical potential $\mu$ and the norm of the superfluid velocities $v_{I}^{s}=|\bm{v^{s}_{I}}|$. The local first law of thermodynamics reads
$d\epsilon=$ $Tds+\mu^{I}dn_{I}+\bm{v^{n}}\cdot d\bm{g}$ (4)
$+\bm{h^{I}}\cdot d\bm{v^{s}_{I}}+Nd(\bm{v^{s}_{1}}\cdot\bm{v^{s}_{2}})\,$,
where $s$ is the entropy density, $\bm{v^{n}}$ is the normal fluid velocity, $\bm{h_{I}}=n^{s}_{I}(\bm{v^{s}_{I}}-\bm{v^{n}})$ is the conjugate source to the superfluid velocity and $n^{s}_{I}$ is the superfluid density of the $I^{\textrm{th}}$ component (here, $I$ is not summed over). In contrast to the single component case Chaikin and Lubensky (1995), there is an additional cross term which accounts for the energy change due to relative motion between the two components of superfluids. [^2] In principle in absence of boost invariance there is a relative coefficient between $\bm{v^{s}_{I}}$ and $\bm{v^{n}}$ in $\bm{h_{I}}$ Armas and Have (2024). As it plays no role in our analysis, we have set it to unity. By introducing $\bm{\bar{h}_{I}}=\bm{h_{I}}+N\sigma^{IJ}\bm{v^{s}_{J}}$ with $\sigma^{IJ}=\left(\begin{smallmatrix}0&1\\ 1&0\end{smallmatrix}\right)$, this cross term can be absorbed and one has
@@
We now consider non-ideal corrections to the ideal order constitutive relations, labeled by tildes. One central hypothesis of hydrodynamics is that of local equilibrium, ie we coarse-grain the system over scales larger than the local equilibration scales and define local equilibrium thermodynamic quantities, $\epsilon(t,x^{i})$, $n(t,x^{i})$, etc. The global version of the second law of thermodynamics can then be rewritten locally as the divergence of an entropy current, with $s(t,x^{i})$ and $T(t,x^{i})$ the non-equilibrium entropy density and temperature:
$T\partial_{t}s+T\partial_{i}j^{i}_{s}\equiv\Delta\geq 0\,.$ (6)
$\Delta\geq 0$ is the local version of positivity of entropy production , with $j^{i}_{s}$ the entropy current.
The first law of the thermodynamics can also be promoted to a local version, and used to express $\partial_{t}s$ and $\partial_{i}s$ in terms of the other quantities in (5). After using the evolution equations (2), (3) and doing some integration by parts Chaikin and Lubensky (1995), imposing the absence of entropy production at ideal level gives the constitutive relations for the currents in equilibrium state compatible with parity and time-reversal invariance:
@@
$j^{i}_{s}=sv^{ni}+\tilde{j}^{i}_{s}\,$,
where $p=-\epsilon+sT+n_{I}\mu_{I}+v^{ni}g_{i}$, while tilded quantities are higher-order in derivatives corrections to the ideal constitutive relations.  [^3] While not manifest from (7), the stress tensor $\tau^{ij}$ is symmetric for isotropic states. This can be verified by inverting the relation between $\bm{\bar{h}_{I}^{i}}$ and $\bm{v_{s}^{I}}$, working out the static susceptibilities $\chi_{v_{s}^{Ii}g^{j}}$, using Onsager $\chi_{v_{s}^{Ii}g^{j}}=\chi_{g^{i}v_{s}^{Ij}}$ and deducing the expression for $g^{i}$. Substituting in (7), $\tau^{ij}=\tau^{ji}$. Then, the local second law (6) becomes
$\Delta=-\tilde{j}_{s}^{i}\partial_{i}T/T-\tilde{j}_{I}^{i}\partial_{i}\mu^{I}-\tilde{\tau}^{ji}\partial_{i}v^{n}_{j}-\tilde{\mu^{I}}\partial_{i}\bar{h}_{I}^{i}\,$, (8)
where $\tilde{\mu^{I}}$ are derivative corrections in the Josephson equations. We can now use the second law constraint $\Delta\geq 0$ to determine the most general form of the non-ideal constitutive relations for $\tilde{j}^{i}_{I}$, $\tilde{j}^{i}_{s}$, $\tilde{\mu}_{I}$ and $\tilde{\tau}^{ij}$ compatible with positivity of entropy production. To proceed, a choice of frame must be made, ie a choice of out-of-equilibrium definition of temperature, chemical potentials and velocities, Kovtun (2012). We choose a special ‘thermodynamic’ frame Jensen et al. (2012); Armas and Jain (2021) where only spatial fluxes $\tilde{j}^{i}_{I}$, $\tilde{j}^{i}_{s}$, $\tilde{\mu}_{I}$ and $\tilde{\tau}^{ij}$, not the densities $O_{A}$, receive derivative corrections, and time derivatives are removed by using the ideal EoMs.
II.2 Linear analysis and the instabilities
Next, we turn to the dynamical instability from hydrodynamics. One can linearize the EoMs around equilibrium state by perturbing the thermodynamic quantities $O_{A}=O_{A0}+e^{-i\omega t+i\bm{k}\cdot\bm{x}}\delta O_{A}$ and their sources $s_{A}=s_{A0}+e^{-i\omega t+i\bm{k}\cdot\bm{x}}\delta s_{A}$, where $O_{A}=(s,\bm{g},n_{I},\bm{v^{s}_{I}})$ and $s_{A}=(T,\bm{v^{n}},\mu_{I},\bm{\bar{h}_{I}})$. They are related by the static susceptibility matrix $\bar{\chi}_{AB}=\delta O_{A}/\delta s_{B}=-\delta^{2}\bar{F}/\delta s_{A}\delta s_{B}$, where $\bar{F}=-T\mathrm{ln}Z-\int d^{d}x\,\bm{v^{s}_{I}}\cdot\bm{\bar{h}_{I}}$ is the Legendre-transformed thermal Gibbs free energy and $Z$ static partition function. Given our choice of frame, the linearized EoMs in Fourier space are